1) The total profit for a company in February was 20 percent higher than it was in january. The total profit for the 2 months was $157498. Find the profit for each month.
2)your weekly paycheck is 15 percent less than your coworker's. The total of your paycheck and your coworker's paychek is $645. Find the amount of each person's paychek.

Answer by mananth(12270)

1) The total profit for a company in February was 20 percent higher than it was in january. The total profit for the 2 months was $157498. Find the profit for each month.

Let the profit in january be x
20/100 *x is the increase in February
Profit in february will be 20/100*x+x
The total profit for two months was $ 157498
x+ 20/100*x +x = 157498
2x+1/5 *x =157498
10x+x /5 = 157498
11x=5*157498
x=157498*5/11

x= $71590
157498-71590=85908

January profit = $71590
February profit =$85908
-------------------------------

your weekly paycheck is 15 percent less than your coworker's. The total of your paycheck and your coworker's paychek is $645. Find the amount of each person's paychek.
Let the worker's pay check be x
you get 15/100* x less
so you get x- 15/100 *x
The total pay check is $645
x+( x-15/100 *x)= 645
x+(100x-15 / 100 *x) =645
x+(85x/100)=645
100x+85x /100 =645
185x=64500
x=$348.65 worker's pay check
you will get 15% less than 348.65
$296.35